Required information [The following information applies to the questions displayed below.] On January 1, 2021, the AlleghenyEnd of Period Depreciation for the Period Beginning of Period Book || Depreciation Depreciation Rate (%) | Expense Value Annu3. Units of production (units produced in 2021, 42,000; units produced in 2022, 37,000). (Round Depreciation per unit rate
Answer #1

1) straight line

Choose numerator / Choose denominator = Annual depreciation expense
Depreciable base / Useful life = Annual depreciation expense
(162000-8000) = 154000 / 10 = 15400
Depreciation expense
2021 15400
2022 15400

2) Double decline

Depreciation for the period End of period
Annual period Beginning of period book value Depreciation rate Depreciation expense Accumulated depreciation Ending book value
2021 162000 20% 32400 32400 129600
2022 129600 20% 25920 58320 103680

3) Unit of production

(Original cost-salvage value)/estimated production
(162000-8000/350000) = 0.44
Calculate depreciation expense : 2021
Depreciation per unit rate 0.44
unit produced in 2021 42000
Depreciation in 2021 18480
Calculate 2022 depreciation expense
Depreciation per unit rate 0.44
Unit produced in 2022 37000
Depreciation in 2022 16280